# Enhanced Partially Relevant Video Retrieval through Inter- and Intra-Sample Analysis with Coherence Prediction

### v2

@@ 3.3. Inter Correlation Enhancement
Figure 3. The ICE module employs a two-stage selection process for pseudo label assignment. It first computes the similarity between unpaired video moments and text features in a mini-batch. The pairs with mutual maximum similarity are selected as candidate pairs. Then, only those pairs with a similarity higher than the threshold are retained, ensuring high-confidence pseudo labels for subsequent training.
@@
To achieve this, we propose the Inter Correlation Enhancement (ICE) module. Technically, we aim to additionally learn these unlabeled but correlated visual-text relationships, thus constructing a richer semantic space. We focus on exploiting moment-text pairs that are most likely to be semantically corresponding while filtering out those semantically irrelevant ones. The pseudo label assignment process is shown in Figure 3.
@@
(6) $~\hat{j}=\mathop{\arg\max}\limits_{j\in\{1,\cdots,n\}}(\{s_{\bar{i}j}\}),~\hat{i}=\mathop{\arg\max}\limits_{i\in\{1,\cdots,n_{m}\}}(\{s_{i\bar{j}}\}),$
if $\hat{i}=\bar{i}$ and $\hat{j}=\bar{j}$, they are mutually the most similar to each other and we deem them as a candidate correlated pair.
@@
Finally, we obtain $n_{c}$ correlated pairs through the above two-stage selection. These pairs then form a mini-batch of size $n_{c}$ and we align them using contrastive learning losses defined in Equation (3) and (4). The loss of ICE is denoted as $\mathcal{L}_{ice}$.
3.4. Intra Redundancy Mining
@@
To mitigate the intra-sample redundancy, we propose the Intra Redundancy Mining (IRM) module. It identifies redundant and query-irrelevant moment features by analyzing relations among text, video, and moment features. Unlike conventional negative samples (e.g., randomly sampled unpaired videos), these mined negatives exhibit higher feature similarity to the target moment, presenting a more challenging discrimination task. By forcing the model to distinguish redundant moments and query-relevant moments, we promote the learning of more discriminative and query-relevant video representations, improving retrieval performance.
@@
We construct redundant moment features from two different perspectives. Technically, we model the semantic variation in the latent space (Wu et al., 2021; Qiu et al., 2025; Li et al., 2025) between the key moment feature $m^{k}$, video feature $v$, and query feature $q$ to obtain the video-view redundant feature $r_{v}\in\mathbb{R}^{D}$ and query-view redundant feature $r_{q}\in\mathbb{R}^{D}$:
(7) $r_{v}=FC(v-m^{k}),~r_{q}=FC(v-q),$
@@
where $FC$ is a fully-connected layer to bridge the gap between video-moment and video-query differences. The intuition is that the video-level feature $v$ contains global semantics of all moments within the video. By subtracting the key moment feature $m^{k}$ from the video-level feature $v$, we obtain the video-view redundant feature $r_{v}$, which is deemed to be irrelevant to the text. Similarly, we use a symmetrical operation to obtain the query-view redundant feature $r_{q}$, which is regarded to be irrelevant to the target moment. Since both redundant features $r_{v}$ and $r_{q}$ should be semantically irrelevant to the key moment feature and query feature, we employ $r_{v}$ and $r_{q}$ as negative samples in the moment-level similarity learning (Section 3.2.3), where the loss function is denoted as $\mathcal{L}_{neg}$.
@@
Moreover, we propose to directly align $r_{v}$ and $r_{q}$. As the two redundant features are obtained at the video-view and query-view respectively, their alignment boosts the learning of underlying relationships between videos and text queries in the cross-modal semantic space. The corresponding loss $\mathcal{L}_{red}$ is also the jointly application of Equation (3) and (4). The overall training objective of this module is the sum of the above two losses:
(8) $\mathcal{L}_{irm}=\mathcal{L}_{neg}+\mathcal{L}_{red}.$
3.5. Temporal Coherence Prediction
@@
Figure 4. The pipeline of the Temporal Coherence Prediction (TCP) module. The frame features are first divided into distinct groups, and group labels are assigned accordingly. Then, a subset of frame features is randomly selected and shuffled, and the group labels of these frames are predicted.
@@
While our ICE and IRM modules effectively leverage inter-sample correlations and mitigate intra-sample redundancy, their performance fundamentally depends on the model’s temporal modeling ability. Clearer temporal feature distinctions enable more accurate cross-modal correlation measurement in ICE and more precise redundancy identification in IRM. To further enhance these capabilities, we introduce the Temporal Coherence Prediction (TCP) module, which strengthens temporal structure learning through self-supervised learning. The pipeline is shown in Figure 4.
@@
Specifically, the TCP module incorporates an auxiliary sequence prediction task. To capture the temporal structure of videos, we leverage the continuity between adjacent frames to divide the video sequence into coherent temporal segments. Inspired by HERO (Li et al., 2020), we shuffle the video features to form an unordered sequence. The model is trained to predict the original temporal order. By explicitly modeling temporal coherence, this task helps the model to learn the coherence and causality between video events. It is worth noting that, unlike HERO, which assigns extensive unique sequence labels, TCP adopts grouped labeling. We identify an issue in HERO’s labeling method. Predicting unique labels is overly hard and introduces noise. We address it by grouped labeling, reducing task complexity for robust temporal modeling. We also conduct ablations on group numbers in Section 4.3.3. Note that TCP is applied to both the video-level branch and moment-level branch in Section 3.2.2 to enable multi-grained temporal coherence learning. We mainly take the video-level branch for explanation.
@@
Concretely, we first divide frame features into $g$ groups and each frame feature is given its group label ranging from $1$ to $g$. Every $T_{f}/g$ consecutive frame features are assigned the same group label. Group labels of the original encoded frame features $V_{f}$ are denoted as $y_{o}\in\mathbb{R}^{T_{f}}$. We then randomly select and shuffle some frame features and encode them through the video encoder. Note that shuffling a small or large length of frames might lead to suboptimal performance. Too few shuffled frames provide insufficient learning signals, while excessive shuffling disrupts temporal coherence and makes the task overly challenging. Therefore, we only shuffle 25% frames. The shuffled frame features and their shuffled group labels are represented as $\hat{V_{f}}=\left\{\hat{f_{i}}\right\}_{i=1}^{T_{f}}\in\mathbb{R}^{T_{f}\times D}$ and $y_{s}\in\mathbb{R}^{T_{f}}$, respectively. Then we adopt a classifier to predict the group labels of both the original and shuffled frame features $V_{f}$ and $\hat{V_{f}}$:
(9) $),p_{o}=Softmax(CLS(V_{f})),~p_{s}=Softmax(CLS(\hat{V_{f}}$)),
@@
3.6. Inference
During inference, the ICE, IRM, and TCP modules are deactivated. The similarity of the given video-text pair $\mathcal{V}$ and $\mathcal{T}$ is the weighted sum of moment-level and video-level similarities in Section 3.2.3, following prior works (Dong et al., 2022a; Wang et al., 2024b; Dong et al., 2023; Jiang et al., 2023; Yin et al., 2024):
(11) $\mathcal{S}\left(\mathcal{V},\mathcal{T}\right)=\alpha~\mathcal{S}_{m}\left(\mathcal{V},\mathcal{T}\right)+\left(1-\alpha\right)\mathcal{S}_{v}\left(\mathcal{V},\mathcal{T}\right),$
where $\alpha\in\left[0,1\right]$ is a hyperparameter to balance the similarities.
Table 1. Performance comparison results on TVR, ActivityNet Captions, and Charades-STA. The best results are highlighted in bold and the second-best outcomes are underlined. We achieve state-of-the-art results on all three datasets across all metrics. denotes using additional CLIP-B/32 model for knowledge distillation. indicates using Large Multimodal Models (LMMs).
@@ 3.6. Inference
DE++ (Dong et al., 2021) TPAMI’2021 8.8 21.9 30.2 67.4 128.3 5.3 18.4 29.2 68.0 121.0 1.7 5.6 9.6 37.1 54.1
RIVRL (Dong et al., 2022b) TCSVT’2022 9.4 23.4 32.2 70.6 135.6 5.2 18.0 28.2 66.4 117.8 1.6 5.6 9.4 37.7 54.3
Cap4Video (Wu et al., 2023) CVPR’2023 10.3 26.4 36.8 74.0 147.5 6.3 20.4 30.9 72.6 130.2 1.9 6.7 11.3 45.0 65.0
ReLoCLNet (Zhang et al., 2021b) SIGIR’2021 10.7 28.1 38.1 80.3 157.1 5.7 18.9 30.0 72.0 126.6 1.2 5.4 10.0 45.6 62.3
CONQUER (Hou et al., 2021) MM’2021 11.0 28.9 39.6 81.3 160.8 6.5 20.4 31.8 74.3 133.1 1.8 6.3 10.3 47.5 66.0
MS-SL (Dong et al., 2022a) MM’2022 13.5 32.1 43.4 83.4 172.4 7.1 22.5 34.7 75.8 140.1 1.8 7.1 11.8 47.7 68.4
JSG (Chen et al., 2023) MM’2023 - - - - - 6.8 22.7 34.8 76.1 140.5 2.4 7.7 12.8 49.8 72.7
UMT-L (Li et al., 2023) ICCV’2023 13.7 32.3 43.7 83.7 173.4 6.9 22.6 35.1 76.2 140.8 1.9 7.4 12.1 48.2 69.6
PEAN (Jiang et al., 2023) ICME’2023 13.5 32.8 44.1 83.9 174.2 7.4 23.0 35.5 75.9 141.8 2.7 8.1 13.5 50.3 74.7
GMMFormer (Wang et al., 2024b) AAAI’2024 13.9 33.3 44.5 84.9 176.6 8.3 24.9 36.7 76.1 146.0 2.1 7.8 12.5 50.6 72.9
BGM-Net (Yin et al., 2024) TOMM’2024 14.1 34.7 45.9 85.2 179.9 7.2 23.8 36.0 76.9 143.9 1.9 7.4 12.2 50.1 71.6
PBU (Moon et al., 2025) ICCV’2025 15.4 35.9 47.5 86.4 185.1 7.9 24.9 37.2 77.3 147.4 - - - - -
ARL (Cho et al., 2025) AAAI’2025 15.6 36.3 47.7 86.3 185.9 8.3 24.6 37.4 78.0 148.3 - - - - -
HLFormer (Jun et al., 2025) ICCV’2025 15.7 37.1 48.5 86.4 187.7 8.7 27.1 40.1 79.0 154.9 2.6 8.5 13.7 54.0 78.7
InternVideo2 (Wang et al., 2024a) ECCV’2024 13.8 32.9 44.4 84.2 175.3 7.5 23.4 36.1 76.5 143.5 1.9 7.5 12.3 49.2 70.9
DL-DKD (Dong et al., 2023) ICCV’2023 14.4 34.9 45.8 84.9 179.9 8.0 25.0 37.5 77.1 147.6 - - - - -
MGAKD (Zhang et al., 2025) TOMM’2025 16.0 37.8 49.2 87.5 190.5 7.9 25.7 38.3 77.8 149.6 - - - - -
Ours - 17.5 39.0 49.9 87.6 194.0 10.1 28.6 41.9 79.8 160.4 2.9 9.2 14.9 54.3 81.3
4. Experiments
4.1. Experimental Setup
4.1.1. Datasets
@@
To evaluate the effectiveness of our method on the PRVR task, we utilize three widely-used datasets with untrimmed long videos: TVR (Lei et al., 2020), ActivityNet Captions (Krishna et al., 2017), and Charades-STA (Gao et al., 2017). Note that we do not use moment-level annotations provided by these datasets in the setting of PRVR. Unlike conventional T2VR datasets (Xu et al., 2016; Chen and Dolan, 2011; Wang et al., 2019) using pre-trimmed short clips, PRVR employs untrimmed long videos where the content is only partially relevant to text queries.
TVR is collected from six TV shows, containing 21.8K videos and 109K text queries. Each video has five textual descriptions on average, depicting different moments in the video. Following (Dong et al., 2022a), the training set consists of 17,435 videos with 87,175 moments, while the test set contains 2,179 videos with 10,895 moments.
ActivityNet Captions consists of a wide range of human activities, sourced from 20K YouTube videos. On average, each video has 3.7 moments with sentence descriptions. We adopt the same data partition as in (Dong et al., 2022a, 2023; Wang et al., 2024b), where 10,009 and 4,917 videos (37,421 and 17,505 moments) are utilized as the training and test sets.
Charades-STA primarily comprises videos that depict indoor activities, including 6,670 videos and 16,128 moment annotations. We utilize the official data partition, 12,408 and 3,720 moment-sentence pairs for training and evaluation, respectively.
4.1.2. Evaluation Metrics

### v4

@@ 3.3 Inter Correlation Enhancement
Figure 3: The ICE module employs a two-stage selection process for pseudo label assignment. It first computes the similarity between unpaired video moments and text features in a mini-batch. The pairs with mutual maximum similarity are selected as candidate pairs. Then, only those pairs with a similarity higher than the threshold are retained, ensuring high-confidence pseudo labels.
@@
To achieve this, we propose the Inter Correlation Enhancement (ICE) module. Technically, we aim to learn these unlabeled but correlated visual-text relationships, thus constructing a richer semantic space. We exploit moment-text pairs that are most likely to be semantically corresponding while filtering out those semantically irrelevant ones. The pseudo label assignment process is shown in Figure 3.
@@
$~\hat{j}=\mathop{\arg\max}\limits_{j\in\{1,\cdots,n\}}(s_{\bar{i}j}),~\hat{i}=\mathop{\arg\max}\limits_{i\in\{1,\cdots,n_{m}\}}(s_{i\bar{j}}),$ (4)
if $\hat{i}=\bar{i}$ and $\hat{j}=\bar{j}$, they are mutually the most similar to each other, and we deem them as a candidate pair.
@@
Finally, we obtain $n_{c}$ correlated pairs through the above two-stage selection. These pairs then form a mini-batch of size $n_{c}$, and we align them using contrastive learning losses $\mathcal{L}^{trip}$ and $\mathcal{L}^{nce}$. The loss of ICE is denoted as $\mathcal{L}_{ice}$.
3.4 Intra Redundancy Mining
@@
To mitigate the intra-sample redundancy, we propose the Intra Redundancy Mining (IRM) module. It identifies redundant and query-irrelevant moment features by analyzing relations among text, video, and moment features. Unlike conventional negative samples (e.g., randomly sampled unpaired videos), these mined negatives exhibit higher feature similarity to the target moment, presenting a more challenging discrimination task. By forcing the model to distinguish redundant moments and query-relevant moments, we promote the learning of more discriminative fine-grained video representations, improving retrieval performance.
@@
We construct redundant moment features from two perspectives. Technically, we model semantic variation in latent space between the key moment feature $m_{k}$ (Eq. 2), video feature $v$, and query feature $q$ to obtain video-view and query-view redundant features $r_{v}\in\mathbb{R}^{D}$ and $r_{q}\in\mathbb{R}^{D}$. Since semantic relationships are often captured as linear translations in latent space [mikolov-etal-2013-linguistic, glove] and different modalities are mapped to the shared space with linear structure [10418785], we adopt subtraction to remove specific semantics [TME]:
$r_{v}=FC(v-m_{k}),~r_{q}=FC(v-q),$ (5)
@@
where $FC$ is a fully-connected layer to bridge the gap between video-moment and video-query differences, and $m_{k}$ is obtained in Eq. 2. The intuition is that the video-level feature $v$ contains global semantics of all moments within the video. By subtracting the key moment feature $m_{k}$ from the video-level feature $v$, we obtain the video-view redundant feature $r_{v}$, which is deemed to be irrelevant to the text query. Similarly, we use a symmetrical operation to obtain the query-view redundant feature $r_{q}$, which is regarded as irrelevant to the target moment. Since both redundant features $r_{v}$ and $r_{q}$ should be semantically irrelevant to the key moment feature and query feature, we employ $r_{v}$ and $r_{q}$ as negative samples in the moment-level similarity learning (Section 3.2.3), where the loss function is denoted as $\mathcal{L}_{neg}$.
@@
Moreover, we propose to directly align $r_{v}$ and $r_{q}$. As the two redundant features are obtained at the video-view and query-view respectively, their alignment boosts the learning of underlying relationships between videos and text queries in the cross-modal semantic space. The corresponding loss $\mathcal{L}_{red}$ is also the jointly application of $\mathcal{L}^{trip}$ and $\mathcal{L}^{nce}$. The overall training objective of this module is the sum of the above two losses: $\mathcal{L}_{irm}=\mathcal{L}_{neg}+\mathcal{L}_{red}$.
3.5 Temporal Coherence Prediction
@@
Figure 4: The pipeline of the Temporal Coherence Prediction (TCP) module. The frame features are first divided into distinct groups. Then, a subset of frame features is randomly selected and shuffled, and the group labels of frames are predicted.
@@
While our ICE and IRM modules effectively leverage inter-sample correlations and mitigate intra-sample redundancy, their performance fundamentally depends on the model’s temporal modeling ability. Clearer temporal feature distinctions enable more accurate cross-modal correlation measurement in ICE and more precise redundancy identification in IRM. To further enhance these capabilities, we introduce the Temporal Coherence Prediction (TCP) module, which strengthens discrimination of fine-grained moment-level semantics through self-supervised learning. Ablation studies in Sec. 4.3.1 indicate that TCP further enhances ICE and IRM. The pipeline of TCP is shown in Figure 4.
@@
To capture the temporal structure of videos, we leverage the continuity between adjacent frames to divide the video sequence into coherent temporal segments. Inspired by Jigsaw Puzzle Solving in image representation learning [wei2019iterative, jing2020self, zhuge2021kaleido], we shuffle the video features to form an unordered sequence. The model is trained to predict the original temporal order. By explicitly modeling temporal coherence, this task helps the model to learn the coherence and causality between video events. Instead of assigning extensive unique sequence labels, we adopt grouped labeling. Predicting unique labels in a long video sequence is overly hard and introduces noise. We address it by grouped labeling, reducing task complexity for robust temporal modeling. We also conduct ablation studies on group numbers in Section 4.3.3. Note that TCP is applied to both the video-level branch and moment-level branch in Section 3.2.2 to enable multi-grained temporal coherence learning. We mainly take the video-level branch for explanation.
@@
Concretely, we first divide frame features into $g$ groups, and each frame feature is given its group label ranging from $1$ to $g$. Every $T_{f}/g$ consecutive frame features are assigned the same group label. Group labels of the original encoded frame features $V_{f}$ are denoted as $y_{o}\in\mathbb{R}^{T_{f}}$. We then randomly select and shuffle some frame features and encode them through the video encoder. Note that shuffling a small or large length of frames might lead to suboptimal performance. Too few shuffled frames provide insufficient learning signals, while excessive shuffling disrupts temporal coherence and makes the task overly challenging. Therefore, we only shuffle 25% frames. The shuffled frame features and their shuffled group labels are represented as $\hat{V_{f}}=\left\{\hat{f_{i}}\right\}_{i=1}^{T_{f}}\in\mathbb{R}^{T_{f}\times D}$ and $y_{s}\in\mathbb{R}^{T_{f}}$, respectively. Then we adopt a classifier to predict the group labels of both the original and shuffled frame features $V_{f}$ and $\hat{V_{f}}$:
$),p_{o}=Softmax(CLS(V_{f})),~p_{s}=Softmax(CLS(\hat{V_{f}}$)), (6)
@@
3.6 Inference
During inference, the ICE, IRM, and TCP modules are deactivated. The similarity of the video-text pair $\mathcal{V}$ and $\mathcal{T}$ is the weighted sum of moment-level and video-level similarities in Sec. 3.2.3, following prior works [MSSL, GMMFORMER, DKD, PEAN, yin2024exploiting]:
$\mathcal{S}\left(\mathcal{V},\mathcal{T}\right)=\alpha~\mathcal{S}_{m}\left(\mathcal{V},\mathcal{T}\right)+\left(1-\alpha\right)\mathcal{S}_{v}\left(\mathcal{V},\mathcal{T}\right),$ (8)
where $\alpha\in\left[0,1\right]$ is a hyperparameter to balance similarities.
Table 1: Performance comparison results on TVR, ActivityNet Captions, and Charades-STA. The best results are highlighted in bold and the second-best outcomes are underlined. We achieve state-of-the-art results on all three datasets across all metrics. * means built upon MS-SL [MSSL]. denotes using additional CLIP-B/32 model. indicates using Large Multimodal Models (LMMs).
@@ 3.6 Inference
Cap4Video [wu2023cap4video] CVPR’2023 10.3 26.4 36.8 74.0 147.5 6.3 20.4 30.9 72.6 130.2 1.9 6.7 11.3 45.0 65.0
JSG [JSG] MM’2023 - - - - - 6.8 22.7 34.8 76.1 140.5 2.4 7.7 12.8 49.8 72.7
PEAN [PEAN] ICME’2023 13.5 32.8 44.1 83.9 174.2 7.4 23.0 35.5 75.9 141.8 2.7 8.1 13.5 50.3 74.7
UMT-L [li2023unmasked] ICCV’2023 13.7 32.3 43.7 83.7 173.4 6.9 22.6 35.1 76.2 140.8 1.9 7.4 12.1 48.2 69.6
InternVideo2 [wang2024internvideo2] ECCV’2024 13.8 32.9 44.4 84.2 175.3 7.5 23.4 36.1 76.5 143.5 1.9 7.5 12.3 49.2 70.9
BGM-Net [yin2024exploiting] TOMM’2024 14.1 34.7 45.9 85.2 179.9 7.2 23.8 36.0 76.9 143.9 1.9 7.4 12.2 50.1 71.6
DL-DKD [DKD] ICCV’2023 14.4 34.9 45.8 84.9 179.9 8.0 25.0 37.5 77.1 147.6 - - - - -
RAL* [RAL] EMNLP’2025 14.5 34.3 45.8 84.5 179.1 7.4 23.4 35.4 76.7 143.0 - - - - -
PBU [PBU] ICCV’2025 15.4 35.9 47.5 86.4 185.1 7.9 24.9 37.2 77.3 147.4 - - - - -
SDM* [SDM] AAAI’2025 15.7 36.7 47.9 85.8 186.2 7.2 23.5 35.8 76.9 143.4 - - - - -
HLFormer [HLFormer] ICCV’2025 15.7 37.1 48.5 86.4 187.7 8.7 27.1 40.1 79.0 154.9 2.6 8.5 13.7 54.0 78.7
MGAKD [MGAKD] TOMM’2025 16.0 37.8 49.2 87.5 190.5 7.9 25.7 38.3 77.8 149.6 - - - - -
MS-SL++ [MS-SL++] TPAMI’2025 13.6 33.1 44.2 83.5 174.5 7.0 23.1 35.2 75.8 141.1 1.8 7.6 12.0 48.4 69.7
MS-SL [MSSL] MM’2022 13.5 32.1 43.4 83.4 172.4 7.1 22.5 34.7 75.8 140.1 1.8 7.1 11.8 47.7 68.4
+ Ours - 17.1 37.9 48.2 86.2 189.4 9.8 28.1 41.1 78.6 157.6 2.4 9.1 14.6 53.9 80.0
GMMFormer [GMMFORMER] AAAI’2024 13.9 33.3 44.5 84.9 176.6 8.3 24.9 36.7 76.1 146.0 2.1 7.8 12.5 50.6 72.9
+ Ours - 18.6 40.3 51.1 89.3 199.3 11.5 30.2 43.8 81.6 167.1 3.0 9.5 15.3 55.1 82.9
Base Model - 13.8 32.2 44.1 83.7 173.8 7.0 22.9 35.1 76.3 141.3 1.5 6.8 11.3 49.5 69.1
+ Ours - 17.5 39.0 49.9 87.6 194.0 10.1 28.6 41.9 79.8 160.4 2.9 9.2 14.9 54.3 81.3
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets
@@
To evaluate the effectiveness of our method, we utilize three widely-used datasets with untrimmed long videos: TVR [tvr], ActivityNet Captions [activitynet], and Charades-STA [sta]. Note that we do not use moment-level annotations provided by these datasets in the setting of PRVR. Unlike conventional T2VR datasets [xu2016msr, chen2011collecting, wang2019vatex] using pre-trimmed short clips, PRVR employs untrimmed long videos where the content is only partially relevant to text queries. TVR is collected from six TV shows, containing 21.8K videos and 109K text queries. Each video has five textual descriptions on average, depicting different moments in the video. Following [MSSL], the training set consists of 17,435 videos with 87,175 moments, while the test set contains 2,179 videos with 10,895 moments. ActivityNet Captions consists of a wide range of human activities, sourced from 20K YouTube videos. On average, each video has 3.7 moments with sentence descriptions. We adopt the same data partition as in [MSSL, DKD, GMMFORMER], where 10,009 and 4,917 videos (37,421 and 17,505 moments) are utilized as the training and test sets. Charades-STA primarily comprises videos that depict indoor activities, including 6,670 videos and 16,128 moment annotations. We adopt the official data partition, where 12,408 and 3,720 moment-sentence pairs are used for training and evaluation, respectively.
4.1.2 Evaluation Metrics